# The Multilingual Mind: A Survey of Multilingual Reasoning in Language Models

### v1

@@ 4 Methods
Multilingual reasoning within language models has garnered significant attention in recent years, leading to the development of diverse techniques for enhancing their capabilities across diverse languages. Prior works have explored various directions to improve multilingual reasoning. Building upon this body of work (see Figure 6), we identify four primary thrusts, viz. representation alignment, fine-tuning, prompting, and model editing, that collectively contribute to advancing multilingual reasoning in language models.
@@
a) Representation Alignment. Multilingual reasoning requires consistent representations across languages, but LLMs often struggle due to imbalanced training data. Representation alignment ensures that equivalent concepts share similar embeddings, reducing inconsistencies in cross-lingual inference, essential for reasoning and multilingual generalization. Li et al. (2024b) employs Multilingual Contrastive Learning to align multilingual sentence representations by treating translation pairs as positive samples and pulling their embeddings closer, bridging representation gaps between languages, enhancing cross-lingual reasoning and generation capabilities. Multilingual Alignment Learning is another form of representation alignment that ensures semantic consistency across languages by aligning their representations for improved multilingual performance, where prior works leverage external multilingual models to enhance reasoning by aligning multilingual capabilities with LLMs (Huang et al., 2024b), bridge multilingual encoders with LLMs using minimal parameters to achieve effective alignment without supervision (Yoon et al., 2024), and evaluate the semantic similarity between English and other languages within LLM embeddings to measure alignment quality (Kargaran et al., 2024). Furthermore, an exciting new direction in representation alignment is Multilingual Compositional Learning which constructs compositional representations by combining equivalent token embeddings across multiple languages (Arora et al., 2024).
@@
b) Finetuning. It leverages cross-lingual data and tasks to fine-tune models for enhanced reasoning and comprehension, leading to numerous innovative approaches. One such approach, exemplified by LinguaLIFT (Zhang et al., 2024), uses code-switched fine-tuning along with language alignment layers to effectively bridge the gap between English and languages with fewer resources, helping maintain the nuance and context across linguistic boundaries. Similarly, QuestionAlign (Zhu et al., 2024) takes a step further by aligning questions and responses in multiple languages, thereby enhancing cross-lingual understanding and consistency in reasoning. While these methods have leaned towards extensive fine-tuning, SLAM (Fan et al., 2025) introduces a more parameter-efficient strategy, which selectively tunes only layers critical for multilingual comprehension, significantly lowering the computational demands while still maintaining or even enhancing the model’s reasoning capabilities. Translation has also been harnessed as a powerful tool for knowledge transfer in multilingual settings. TransLLM (Geng et al., 2024), for instance, focuses on translation-aware fine-tuning to align different languages, thereby improving the model’s ability to perform reasoning across languages. This method not only enhances language understanding but also adapts the model for various cross-lingual tasks. For those aiming at more complex reasoning tasks, reasoning-focused fine-tuning has proven beneficial. The Multilingual Chain-of-Thought (mCoT) Instruction Tuning method (Lai and Nissim, 2024) utilizes a dataset specifically curated for reasoning across languages (mCoT-MATH) and combines CoT reasoning with instruction tuning to boost consistency and logical problem-solving in multiple languages. In addition, preference-based techniques to align reasoning outputs across languages emphasizes the use of language imbalance as a reward signal in models like DPO (Direct Preference Optimization) and PPO (Proximal Policy Optimization) (She et al., 2024). Finally, an interesting direction moving forward is the use of curriculum-based and retriever-based fine-tuning techniques to enhance multilingual reasoning (Anand et al., 2024; Bajpai and Chakraborty, 2024).
@@
c) Prompting. Prompting has emerged as a key technique for enhancing how LLMs adapt and reason across different languages. By guiding the model through specific strategies, prompting not only facilitates dynamic adaptation but also addresses the challenges posed by data imbalances, thereby enhancing cross-lingual consistency, logical alignment, and the robustness of reasoning. One effective method is Direct Multilingual Input Prompting, as explored in the study by Sakai et al. (2024). Here, the model directly processes inputs in various native languages without any need for translation, preserving the original linguistic nuances. This approach was notably applied in the paper “Do Moral Judgements” by Khandelwal et al. (2024), where moral scenarios were presented in their native languages to assess the model’s reasoning capabilities directly. Another strategy, Translation-Based Prompting (Liu et al., 2024) uses translation to convert multilingual inputs into a pivot language for processing, where tasks are translated into English for reasoning and translated back to the target language for evaluation (Wang et al., 2024; Zhao and Zhang, 2024), where one can also generate diverse CoT with Negative Rationales (d-CoT-nR) by incorporating both correct and incorrect reasoning paths to refine multilingual reasoning capabilities (Payoungkhamdee et al., 2024).
@@
d) Model Editing. Model (or knowledge) editing is a growing and exciting research area that aims to modify/update the information stored in a model. Formally, model editing strategies update pre-trained models for specific input-output pairs without retraining them and impacting the baseline model performance on other inputs. Multilingual Precision Editing involves making highly specific updates to model knowledge while ensuring minimal impact on unrelated information. Multilingual knowledge Editing with neuron-Masked Low-Rank Adaptation (MEMLA) (Xie et al., 2024) enhances multilingual reasoning by leveraging neuron-masked LoRA-based edits to integrate knowledge across languages and improve multi-hop reasoning capabilities. Further, Multilingual Translation Post-editing refines translations by correcting errors in multilingual outputs for better alignment and fluency, where we can enhance multilingual reasoning by incorporating auxiliary translations into the post-editing process, enabling LLMs to improve semantic alignment and translation quality across languages (Lim et al., 2024).
@@ 5.1 Metrics
Here, we categorize the key evaluation metrics for multilingual reasoning, along with their formal definitions:
1) Accuracy-Based Metrics. These metrics assess overall correctness in reasoning and multilingual benchmarks: i) General Accuracy measures the proportion of correct outputs over total samples and ii) Zero-Shot Accuracy, which evaluates model performance on unseen tasks or categories without fine-tuning.
2) Reasoning and Consistency Metrics. These metrics evaluate logical inference and multi-step reasoning ability: i) Reasoning Accuracy: Assesses correctness in logical and step-by-step reasoning tasks and ii) Path Consistency: Measures coherence between reasoning steps in CoT prompting.
3) Translation and Cross-Lingual Metrics. To ensure multilingual reasoning consistency, models must preserve meaning across languages: i) Translation Success Rate (TSR): Measures correctness and semantic preservation in multilingual translations as the ratio of accurate translations and total translations and ii) Cross-Lingual Consistency: Evaluates whether logically equivalent statements yield consistent reasoning outputs across different languages.
4) Perplexity and Alignment Metrics. These metrics quantify semantic alignment and measure whether embeddings across languages remain consistent: i) Perplexity-Based Alignment ($P_{\text{align}}$):
@@
where $P(x_{i})$ is the model’s probability of predicting token $x_{i}$. Lower perplexity indicates better alignment and ii) Semantic Alignment Score ($S_{\text{align}}$) that measures the cosine similarity between multilingual sentence embeddings:
$S_{\text{align}}=\frac{E_{l}\cdot E_{t}}{\|E_{l}\|\|E_{t}\|},$ (2)
where $E_{l}$ and $E_{t}$ are sentence embeddings in different languages.
@@ 5.2 Performance on Benchmarks
Here, we discuss the performance of the aforementioned methods on standard mathematical reasoning benchmarks (MGSM (Shi et al., 2022) and MSVAMP (Chen et al., 2023)), common sense reasoning (xCSQA (Lin et al., 2021)), and logical reasoning (xNLI (Conneau et al., 2018)) [^1] We only cover benchmarks analyzed by more than four papers..
Figure 7: Accuracy trends of various methodologies on multilingual reasoning benchmarks, including MGSM, MSVAMP, XNLI, and XCSQA. The x-axis represents the date of paper submission on arXiv, while the y-axis indicates percentage accuracy.
In Fig. 7, the x-axis is the timestamp of their arXiv submission, and the y-axis is the predictive performance (accuracy) of the methods on those benchmarks. Next, we describe the four most popular benchmarks and detail the performance of state-of-the-art reasoning techniques on them, highlighting existing gaps in language models that limit their reasoning performance.
MGSM tests multilingual arithmetic reasoning in LLMs with 250 translated math problems in ten diverse languages. While recent trends suggest that advanced post-training techniques like MAPO are crucial for strong performance, fine-tuning strategies may be more impactful than stronger reasoning architectures or relying on the model’s English expertise to improve multilingual performance.
MSVAMP is an out-of-domain multilingual mathematical reasoning dataset comprising 10,000 problems across ten languages and serves as a comprehensive test bed to evaluate LLMs’ generalization in multilingual mathematical contexts. We find that advanced preference optimization (MAPO) achieves much stronger performance than CoT-based fine-tuning, suggesting advanced fine-tuning techniques are a better direction to beat the current best in this benchmark.
xCSQA is a multilingual extension of the CommonsenseQA dataset, encompassing 12,247 multiple-choice questions translated into 15 languages, designed to assess LLMs’ cross-lingual commonsense reasoning capabilities. The current trend in this benchmark shows that stronger fine-tuning strategies like two-step fine-tuning (LingualLIFT) or preference optimization (MAPO) show better performance than selectively fine-tuning specific layers as in SLAM.
xNLI evaluates cross-lingual sentence inference across 15 languages, including low-resource ones. Recent studies suggest that LLM integration with external models (Huang et al., 2024b) and multilingual alignment followed by fine-tuning (Zhang et al., 2024) outperform contrastive learning methods like TCC (Chia et al., 2023), highlighting the need for more structured multilingual adaptation strategies.
@@ 6 Future Directions
As we rapidly develop the new generation of reasoning models, our community must ensure that the models remain unbiased towards languages from underrepresented regions around the world. Looking forward, we call on the community to put their collective efforts into some of the following future directions:
1. Multilingual Alignment and Reasoning Transfer. A key challenge in multilingual reasoning is the lack of data in different languages. One promising solution is to leverage existing large datasets and transfer/distill their knowledge in the representation space, as demonstrated by (Yoon et al., 2024; Huang et al., 2024b). Future research should focus on improving cross-lingual knowledge transfer techniques, enabling models to use high-resource languages as a bridge to enhance reasoning consistency in low-resource languages. Another direction is to generate synthetic datasets using techniques like back-translation and data augmentation, tailored specifically for reasoning tasks.
2. Explainable and Interpretable Reasoning. Ensuring faithful and interpretable reasoning in multilingual LLMs is challenging due to linguistic diversity, translation ambiguities, and reasoning inconsistencies. Studies on English CoT reasoning (Tanneru et al., 2024; Lobo et al., 2024) highlight faithfulness issues, which become more severe when extended to low-resource languages. Causal reasoning can enhance cross-lingual alignment, improving interpretability by uncovering cause-effect relationships across languages. Future research should focus on integrating causal reasoning and multilingual CoT frameworks to ensure logical coherence, transparency, and trust in multilingual AI systems.
3. Advanced Training and Inference Techniques. While recent advancements in multilingual reasoning have introduced reasoning-aware fine-tuning and multilingual preference optimization techniques, further efforts are needed to improve training paradigms. Notably, Wu et al. (2024) demonstrates that reward signals can be effectively transferred across languages, paving the way for post-training RL methods that improve reasoning in low-resource languages. Additionally, efficient inference-time scaling and agentic frameworks remain under-explored, despite emerging techniques (Khanov et al., 2024; Chakraborty et al., 2024) gaining traction and multi-agent frameworks (Guo et al., 2024) enabling LLMs to simulate agent interactions or refine their reasoning by learning from self-generated reasoning paths.
@@
5. Benchmarks: As multilingual reasoning advances, robust evaluation benchmarks are essential. As reasoning is highly domain-specific in nature, developing targeted benchmarks is crucial, especially in high-stakes fields like healthcare, law, and finance, where accuracy directly affects decision-making.
6. Multimodal Multilingual Reasoning. While there are a few works on visual reasoning in the multilingual context (Das et al., 2024), multimodal reasoning (integrating tables, text, image, audio and video) remains largely unexplored. Advancing this area could enable models handle complex tasks in low-resource languages and incorporate cross-modal reasoning.

### v2

@@ 4 Methods
Multilingual reasoning within LMs has garnered significant attention in recent years, leading to the development of diverse techniques for enhancing their capabilities across diverse languages. Prior works have explored various directions to improve multilingual reasoning. Building upon this body of work (see Fig. 5), we identify four primary thrusts, viz. representation alignment, fine-tuning, prompting, and model editing, collectively contributing to advancing multilingual reasoning in LMs.
@@
a) Representation Alignment. Multilingual reasoning requires consistent representations across languages, but LMs often struggle due to imbalanced training data. Representation alignment ensures that equivalent concepts share similar embeddings, reducing inconsistencies in multilingual inference, vital for reasoning and multilingual generalization. Li et al. (2024b) employs contrastive learning to align multilingual sentence representations by treating translation pairs as positive samples and pulling their embeddings closer, bridging language representation gaps and enhancing model’s cross-lingual reasoning and generation capabilities. Multilingual Alignment Learning is another technique that ensures semantic consistency across languages by aligning their representations for improved multilingual performance (Huang et al., 2024b), bridging multilingual encoders with LLMs using minimal parameters to achieve effective alignment without supervision (Yoon et al., 2024; Kargaran et al., 2024). Similarly, Ruan et al. (2025) integrates all encoder layer representations and employs adaptive fusion-enhanced attention to enable layer-wise alignment between the LLM and multilingual encoder, ensuring consistent cross-lingual representations and improving the model’s multilingual reasoning capabilities. Finally, an exciting new direction is multilingual compositional learning, which constructs compositional representations by combining equivalent token embeddings across multiple languages (Arora et al., 2024) and formalizing problems in an abstract space and solving them step-by-step using self-training for improved alignment across languages (Ranaldi and Pucci, 2025).
@@
b) Finetuning. It leverages cross-lingual data and tasks to fine-tune models for enhanced reasoning and comprehension, leading to numerous innovative approaches. For instance, LinguaLIFT (Zhang et al., 2024a) uses code-switched fine-tuning along with language alignment layers to effectively bridge the gap between English and low-resource languages, helping maintain the nuance and context across linguistic boundaries. Similarly, QuestionAlign (Zhu et al., 2024b) aligns questions and responses in multiple languages, thereby enhancing cross-lingual understanding and consistency in reasoning and Ko et al. (2025) introduces a strategic fine-tuning approach that anchors reasoning in English and then translates results, significantly reducing cross-lingual performance gaps. Strategic fine-tuning using a small but high-quality bilingual dataset can enhance both the reasoning capabilities and non-English language proficiency of LLMs (Ha, 2025). While these methods have leaned towards extensive fine-tuning, SLAM (Fan et al., 2025) introduces a more parameter-efficient strategy and selectively tunes layers critical for multilingual comprehension, significantly lowering the computational demands while still maintaining or even enhancing the model’s reasoning capabilities. Translation has also been harnessed as a powerful tool for knowledge transfer in multilingual settings, where TransLLM (Geng et al., 2024) focuses on translation-aware fine-tuning to align different languages, enhancing language understanding but also adapting the model for various cross-lingual tasks. For those aiming at more complex reasoning tasks, reasoning-focused fine-tuning has proven beneficial. The Multilingual CoT (mCoT) instruction tuning method (Lai and Nissim, 2024) utilizes a dataset specifically curated for reasoning across languages and combines CoT reasoning with instruction tuning to boost consistency and logical problem-solving in multiple languages. In addition, preference-based techniques to align reasoning outputs across languages emphasize the use of language imbalance as a reward signal in models like Direct Preference and Proximal Policy Optimization (She et al., 2024). Recent research has demonstrated that Process Reward Modeling offers fine-grained feedback at each step of the reasoning process, only Wang et al. (2025) has shown its application on non-English language. Finally, an interesting direction moving forward is curriculum-based and retriever-based fine-tuning techniques to enhance multilingual reasoning (Anand et al., 2024; Bajpai and Chakraborty, 2024), where models must not only retrieve relevant information but also compare them to evaluate relationships between them (Agrawal et al., 2024; Ranaldi et al., 2025b; Shao et al., 2024; Yang et al., 2025).
@@
c) Prompting. Prompting has emerged as a key technique for enhancing how LLMs adapt and reason across different languages. By guiding the model through specific strategies, prompting facilitates dynamic language adaptation and addresses the data imbalance challenge, thereby enhancing cross-lingual consistency, logical alignment, and the robustness of reasoning. For instance, an effective method is Direct Multilingual Input Prompting (Sakai et al., 2024), where the model directly processes inputs in various native languages without translation, preserving the original linguistic nuances. This approach was notably applied in the paper “Do Moral Judgements” (Khandelwal et al., 2024), where moral scenarios were directly presented in their native languages to assess the model’s reasoning capabilities. Another strategy, Translation-based prompting (Liu et al., 2024) uses translation to convert multilingual inputs into a target language for processing, where tasks are translated into English for reasoning and translated back to the target language for evaluation (Wang et al., 2024; Zhao and Zhang, 2024b). This is also used to generate diverse CoT with Negative Rationales by incorporating both correct and incorrect reasoning paths to refine multilingual reasoning capabilities (Payoungkhamdee et al., 2024). While in-context learning with natural language can be ambiguous and less effective in low-resource languages, program-based demonstrations offer clearer, structured reasoning that transfers better across languages (Ranaldi et al., 2025a). In addition to the above strategies, Dictionary Insertion Prompting (DIP) offers a lightweight and practical alternative by inserting English translations of keywords into non-English prompts, bridging linguistic gaps without full translation and enabling clearer reasoning and improved performance in multilingual tasks (Lu et al., 2024).
Figure 6: Accuracy trends of various methods on multilingual reasoning benchmarks, including MGSM, MSVAMP, XNLI, and XCSQA. The $x$-axis represents the arXiv paper submission date, and the $y$-axis indicates percentage accuracy.
@@
d) Model Editing. Model editing is a growing and exciting research area that aims to modify/update the information stored in a model. Formally, model editing strategies update pre-trained models for specific input-output pairs without retraining them and impacting the baseline model performance on other inputs. Multilingual Precision Editing involves making updates to model knowledge while ensuring minimal impact on unrelated information. Multilingual knowledge Editing with neuron-Masked Low-Rank Adaptation (MEMLA) (Xie et al., 2024) enhances multilingual reasoning by leveraging neuron-masked LoRA-based edits to integrate knowledge across languages and improve multi-hop reasoning capabilities. Further, Multilingual Translation Post-editing refines translations by correcting errors in multilingual outputs for better alignment, where we can enhance multilingual reasoning by incorporating auxiliary translations into the post-editing process, enabling LLMs to improve semantic alignment and translation quality across languages (Lim et al., 2024).
An emerging complementary direction investigates inference-time (test-time) compute scaling in enhancing multilingual reasoning. Recent work shows that scaling up compute for English-centric reasoning language models (RLMs) can significantly improve performance across many languages, including low-resource ones, even surpassing larger models (Yong et al., 2025). While most test-time techniques, such as CoT prompting with trial and error, have primarily focused on English, methods like English-Pivoted CoT training (Tran et al., 2025) exploit the model’s strong English reasoning capabilities to support multilingual tasks, offering a promising path to bridge alignment gaps for underrepresented languages.
@@ 5.1 Metrics
Here, we detail key metrics for evaluating multilingual reasoning, along with their formal definitions:
1) Accuracy. These metrics assess overall correctness in reasoning and multilingual benchmarks: i) General Accuracy measures the proportion of correct outputs over total samples, and ii) Zero-Shot Accuracy, which evaluates model performance on unseen tasks or categories without fine-tuning.
2) Reasoning and Consistency. These metrics evaluate logical inference and multi-step reasoning ability: i) Reasoning Accuracy assesses correctness in logical and step-by-step reasoning tasks and ii) Path Consistency measures coherence between reasoning steps in CoT prompting.
3) Translation and Cross-Lingual. To ensure multilingual reasoning consistency, models must preserve meaning across languages: i) Translation Success Rate measures correctness and semantic preservation in multilingual translations as the ratio of accurate translations and total translations and ii) Cross-Lingual Consistency evaluates whether logically equivalent statements yield consistent reasoning outputs across different languages.
4) Perplexity and Alignment. They quantify semantic alignment and measure whether embeddings across languages remain consistent: i) Perplexity-Based Alignment ($P_{\text{align}}$)
@@
where $P(x_{i})$ is the model’s probability of predicting token $x_{i}$ (lower perplexity means better alignment) and ii) Semantic Alignment measures the cosine similarity between multilingual sentence embeddings: $S_{\text{align}}=\frac{E_{l}\cdot E_{t}}{\|E_{l}\|\|E_{t}\|},$ where $E_{l}$ and $E_{t}$ are sentence embeddings in different languages.
@@ 5.2 Performance on Benchmarks
Here, we discuss the performance of the aforementioned methods on standard mathematical (MGSM (Shi et al., 2022), MSVAMP (Chen et al., 2023)), commonsense (xCSQA (Lin et al., 2021)), logical (xNLI (Conneau et al., 2018)) reasoning benchmarks [^1] We only cover benchmarks analyzed by more than four papers.. Next, we describe the four most popular benchmarks and detail the performance of reasoning techniques, highlighting existing model gaps that limit their reasoning performance.
MGSM tests multilingual arithmetic reasoning in LMs with 250 translated math problems in ten diverse languages. While recent trends suggest that advanced post-training techniques like MAPO are key for strong performance, fine-tuning strategies may be more impactful than stronger reasoning architectures or relying on the model’s English expertise to improve multilingual performance.
MSVAMP is an out-of-domain multilingual mathematical reasoning dataset comprising 10k problems across ten languages and serves as a comprehensive test bed to evaluate LMs’ generalization in multilingual mathematical contexts. We find that advanced preference optimization achieves much stronger performance than CoT-based fine-tuning, suggesting advanced fine-tuning techniques are a better direction to beat the current best in this benchmark.
xCSQA is a multilingual extension of the CommonsenseQA dataset, encompassing 12,247 multiple-choice questions translated into 15 languages, designed to assess LMs’ cross-lingual commonsense reasoning capabilities. The current trend shows that stronger fine-tuning strategies like two-step fine-tuning or preference optimization show better performance than selectively fine-tuning specific layers as in SLAM.
xNLI evaluates cross-lingual inference across 15 languages. Recent studies suggest that LM integration with external models (Huang et al., 2024b) and multilingual alignment followed by fine-tuning (Zhang et al., 2024a) outperform contrastive learning methods like TCC (Chia et al., 2023), highlighting the need for more structured multilingual adaptation strategies.
@@ 6 Future Directions
With the rapid development of reasoning models, our community must ensure that models remain unbiased towards low-resource languages. Looking forward, we call on the community to put their collective efforts into the following directions:
1. Multilingual Alignment and Reasoning Transfer. A key challenge in multilingual reasoning is the lack of data in different languages. One promising solution is to leverage existing large datasets and transfer/distill their knowledge in the representation space (Yoon et al., 2024; Huang et al., 2024b). Future research should develop cross-lingual knowledge transfer techniques, enabling models to use high-resource languages as a bridge to enhance reasoning in low-resource languages. Another direction is to generate synthetic datasets using techniques like back-translation and data augmentation, tailored specifically for reasoning tasks.
2. Explainable and Interpretable Reasoning. Ensuring faithful reasoning in multilingual LLMs is challenging due to linguistic diversity, translation ambiguities, and reasoning inconsistencies. Studies on English CoT reasoning  (Tanneru et al., 2024; Lobo et al., 2024) highlight faithfulness issues, which become more severe when extended to low-resource languages. Causal reasoning can enhance cross-lingual alignment, improving interpretability by uncovering cause-and-effect relationships across languages. Future research should focus on integrating causal reasoning and multilingual CoT frameworks to ensure logical coherence, transparency, and trust in multilingual AI systems.
3. Advanced Training and Inference Techniques. While recent advancements in multilingual reasoning have introduced reasoning-aware fine-tuning and multilingual preference optimization techniques, further efforts are needed to improve training paradigms. Some exciting techniques in this direction includes post-training RL methods that improve reasoning in low-resource languages (Wu et al., 2024) and efficient inference-time scaling and Agentic frameworks (Khanov et al., 2024; Chakraborty et al., 2024). Preliminary post-training works (Xuan et al., 2025) show that they yield mixed results across languages, with effectiveness depending on the base model and required degree of linguistic diversity, highlighting the need for language inclusive training approaches.
@@
5. Multimodal Multilingual Reasoning. While there are a few works on visual reasoning in the multilingual context (Das et al., 2024; Gao et al., 2025; Ghosh et al., 2024c), multimodal reasoning (integrating tables, text, image, audio, and video) remains largely unexplored. Advancing this area could enable models to handle complex tasks in low-resource languages and incorporate cross-modal reasoning.
6.New Benchmarks: As multilingual reasoning advances, robust evaluation benchmarks are essential because reasoning is highly domain-specific in nature, developing targeted benchmarks is crucial, especially in high-stakes fields like healthcare, law, and finance, where accuracy directly affects decision-making. For instance, Xue et al. (2024) introduces FAMMA which shows significant challenges in the field of Financial Question Answering.
7. Efficient Reasoning Models. An emerging direction in reasoning research is enhancing resource efficiency in reasoning-aware models. Recent works like Ning et al. (2024) propose strategies for more efficient reasoning, reducing computational costs while maintaining logical consistency. However, this area remains largely unexplored in multilingual settings, offering a key opportunity to develop scalable reasoning models that generalize across languages with minimal resources.
8. Miscellaneous Tasks. LLMs have achieved remarkable performance across a wide range of tasks; however, they continue to struggle with complex compositional reasoning (Zhao and Zhang, 2024a), often performing only marginally better than random guessing. They also face difficulties in reasoning over longer contexts, particularly in low-resource languages (Hengle et al., 2025). Moreover, their reasoning traces frequently exhibit hallucinations (Sahoo et al., 2024), with models failing to reliably integrate information or recognize missing pieces even when the relevant facts are retrievable.